need help with the addition method
x-2y= -24 you have to subsitute the numbers in the equation. Thanks!
-6x-3y=9

Answer by checkley75(3666)

multiply the first equation by 6 & we get
6x-12y=-144 now add thiese 2 equations
-6x-3y=9
-------------------------------
-15y=-135
y=-135/-15
y=9 answer.
x-2*9=-24
x-18=-24
x=-24+18
x=-6 answer.